OpenStudy (anonymous):

find all real solutions for x for the given equation (4)/(x^4)-(9)/(x^2)+2=0

OpenStudy (saifoo.khan):

\[\frac{4}{x^4} - \frac{9}{x^2} +2 = 0\]Take x^4 \[\frac{4 - 9x^2}{x^4} = -2\] \[4 - 9x^2 = -2x^2 \]

OpenStudy (saifoo.khan):

that's -2x^4.

OpenStudy (saifoo.khan):

\[2x^4 - 9x^2 + 4 = 0\]Now suppose, x^2 = y \[2y^2 - 9y + 4 =0\]

OpenStudy (saifoo.khan):

Can you solve now?